OpenStudy (nbarrera):

Solve. 12 – x = 4 –8 8 10 14

OpenStudy (anonymous):

x = 8

OpenStudy (anonymous):

It becomes 8-x = 0, Thus it's X = 8

OpenStudy (nbarrera):

thx

OpenStudy (anonymous):

@kobe9115 the way described it is new for me, and makes no sense. sorry if i'm incorrect , but wouldn't x=8 because 12-4=8?

OpenStudy (anonymous):

it's the same thing, some people pref to set eqs like this equal to 0 12 - x = 4 subtract 4 from sides 8 - x = 0 add x to both sides 8 = x